# Optimal Methods for Unknown Piecewise Smooth Problems I: Convex Optimization

### v1

@@ 6 Restarted APEX for unknown $\mu^{*}$
By Lemma 6.2, Algorithm 6 will generate a point set $\mathcal{P}_{s}$ as $\left(\sqrt{2\Delta_{s}/\mu_{s}},\sqrt{\mu_{s}\Delta_{s}/2}\right)$-W-certificate for point $\bar{y}^{s-1}$ in function $f$. By Lemma 4.2, the W-certificate and $\mu^{*}$-strong convexity implies that $f\left(\bar{y}^{s-1}\right)-f^{*}\leq\max\left\{1,\mu_{s}/\mu^{*}\right\}\Delta_{s}$. To make sure $f\left(\bar{y}^{s-1}\right)-f^{*}\leq\varepsilon$, we need $\mu_{s}\Delta_{s}\leq\max\left\{\mu_{s},\mu^{*}\right\}\varepsilon$. Notice that Theorem 6.1 is to bound the iteration number for condition $\mu_{s}\Delta_{s}\leq\varepsilon^{2}$. Hence, replacing $\varepsilon$ in the proof procedure of Theorem 6.1 with $\sqrt{\max\left\{\mu_{s},\mu^{*}\right\}\varepsilon}$ and noting that ${\mu}_{s}\geq\frac{1}{4}\mu^{*}$, we have the complexity result . ∎∎
@@ 7 Lower oracle complexity bound for $\mu$-QG and $\left(k,L\right)$-PWS optimization
The proof follows the two-stage approach pioneered by [11, 12]. First, we show that the lower complexity of any first-order deterministic algorithm (Definition 7.3) is lower bounded by that of any zero-respecting algorithm (Definition 7.4) in Proposition 7.1. It thus suffices to prove the lower bound within the zero-respecting algorithm class. Second, inspired by Nesterov’s construction [65], we exhibit a simple $\mu$-strongly convex, $\left(k,L\right)$-PWS function for which every zero-respecting method indeed incurs $\Omega\left(k\sqrt{{L}/{\mu}}\log\left({1}/{\varepsilon}\right)\right)$ queries. Since strongly convex functions are also $\mu$-QG functions, this result also applies to $\mu$-QG functions.
We begin by introducing precise definitions under consideration, including function class (Definition 7.1), orthogonal connection matrix (Definition 7.2), first-order deterministic algorithm (Definition 7.3) and first-order zero-respecting sequences and algorithm (Definition 7.4).
@@
Definition 7.2 (Orthogonal connection matrix).
@@
Definition 7.3 (First-order deterministic algorithm).
@@
$x^{i}=\mathcal{A}\left(\nabla f\left(x^{1}\right),\ldots,\nabla f\left(x^{i-1}\right)\right)\ ,\ \forall i\in\mathbb{N}\ .$ (52)
Definition 7.4 (First-order zero-respecting algorithm).
@@
$\operatorname{supp}\left\{x^{t}\right\}\subseteq\cup_{s<t}\operatorname{supp}\left\{\nabla f\left(x^{s}\right)\right\}\ ,\ \text{for each }t\in\mathbb{N}.$ (53)
@@
$T_{\varepsilon}\left(\mathcal{A}\left[f\right],f\right):=\inf\left\{t\in\mathbb{N}:f\left(x^{t}\right)-f^{*}\leq\varepsilon\right\}\ .$ (54)
@@
$\mathcal{T}_{\varepsilon}\left(\mathbb{A},\mathcal{F}^{d}\right):=\inf_{\mathcal{A}\in\mathbb{A}}\sup_{f\in\mathcal{F}^{d}}T_{\varepsilon}\left(\mathcal{A}\left[f\right],f\right)\ .$ (55)
@@
$\phi_{\mu,L}\left(x\right)=\frac{\mu}{2}\left\|x\right\|^{2}+\frac{L-\mu}{8}x^{\top}Ax-\frac{L-\mu}{4}e_{1}^{\top}x\ ,$ (56)
@@
$\phi_{\mu,L}^{k}\left(x\right)=\max_{i\in\left[k\right]}\left\{\phi_{\mu,L}\left(x_{(i)}\right)\right\}\ ,\ x=\left[x_{(1,:)}^{\top},x_{(2,:)}^{\top},\ldots,x_{(k,:)}^{\top}\right]\in\mathbb{R}^{d}\ ,\ x_{(i,:)}\in\mathbb{R}^{n_{i}}\ ,\ \forall i\in\left[k\right]\ .$ (57)
@@
$\left\|x^{t}-x^{*}\right\|^{2}$ $\geq\left(\frac{\sqrt{L/\mu}-1}{\sqrt{L/\mu}+1}\right)^{2t/k}\left\|x^{0}-x^{*}\right\|^{2}\ ,$ (58)
$f\left(x^{t}\right)-f\left(x^{*}\right)$ $\geq\frac{\mu}{2}\left(\frac{\sqrt{L/\mu}-1}{\sqrt{L/\mu}+1}\right)^{2t/k}\left\|x^{0}-x^{*}\right\|^{2}\ ,$ (59)
@@
It suffices to prove that the claims (58) and (59) hold for any first-order zero-respecting algorithm $\mathcal{A}_{zr}$ by Proposition 7.1. We test $\mathcal{A}_{zr}$ on the PWS instance $f\left(x\right)=\phi_{\mu,L}^{k}\left(x\right)$, where in the definition (57) we set $n_{i}=\infty,i=1,\ldots,k$. For notation simplicity, we denote $q=\frac{\sqrt{L/\mu}-1}{\sqrt{L/\mu}+1}$. Hence, by the fourth property (Item 4) of $\phi_{\mu,L}\left(x\right)$ that $x^{*}=\operatornamewithlimits{argmin}_{x\in\mathbb{R}^{\infty}}f\left(x\right)$ has a block-coordinate structure that is $x^{*}=\left[\left(x_{(1,:)}^{*}\right)^{\top},\left(x_{(2,:)}^{*}\right)^{\top},\ldots,\left(x_{(k,:)}^{*}\right)^{\top}\right]$, where $x_{(j,i)}^{*}=q^{i}$. Without loss of generality, we assume algorithm starts from $x^{0}=0$. Then by construction, we have
$\left\|x^{0}-x^{*}\right\|^{2}=\sum_{j=1}^{k}\sum_{i=1}^{n_{j}}x_{(j,i)}^{*}=\frac{kq^{2}}{1-q^{2}}\ .$ (60)
@@
$\begin{split}&\left\|x^{t}-x^{*}\right\|^{2}\geq\sum_{i=1}^{k}\left\|x_{(i,:)}^{t}-x_{(i,:)}^{*}\right\|^{2}\overset{(a)}{\geq}\sum_{i=1}^{k}\left\|x_{(i,t_{i}+1:\infty)}^{*}\right\|^{2}\\ &\overset{(b)}{=}\sum_{i=1}^{k}\frac{q^{2(t_{i}+1)}}{1-q^{2}}\overset{(c)}{=}\left\|x^{0}-x^{*}\right\|^{2}\sum_{i=1}^{k}\frac{q^{2t_{i}}}{k}\overset{(d)}{\geq}\left\|x^{0}-x^{*}\right\|^{2}q^{2t/k}\end{split}\ ,$ (61)
In the above, $(a)$ holds by discarding the first difference norm between the first $t_{i}$ elements of $x_{(i,:)}^{t}$ and $x_{(i,:)}^{*}$ , $(b)$ follows from $x_{(j,i)}^{*}=q^{i}$, $(c)$ holds by (60) and $(d)$ is because $g:t\to q^{2t}$ is convex, then $\frac{1}{k}\sum_{i=1}^{k}g\left(t_{i}\right)\geq g\left(\frac{1}{k}\sum_{i=1}^{k}t_{i}\right)=q^{2\bar{t}}$. Hence, we complete the first result (58). Combining it and $f\left(x\right)$ is $\mu$-strongly convex, we complete the proof of the second result (59). ∎∎
@@ 8.1 Finite max of convex quadratic functions (MAXQUAD)
Experiment setting. We first evaluate the performance of rAPEX on small-scale MAXQUAD with varying piece number $k$, smoothness $L$ and initial QG estimates $\hat{\mu}_{1}$, as reported in Table 2. In all experiments, we set the dimension ($d$) to $300$, cut number ($m$) to $50$, the optimal QG modulus ($\mu^{*}$) to $1.0$, and terminate rAPEX once rAPEX improves the upper bound (Algorithm 6 enters Line 15) with $\Delta_{s}\leq 10^{-6}$. Ipopt [71] is used as a baseline to provide optimal objective $(f^{*})$.
@@
Figure 5: Absolute error of upper bound and lower bound of rAPEX v.s. Iteration MAXQUAD with $k=50,L=10$ in Table 2.
Figure 6: Absolute error of upper bound and lower bound of rAPEX v.s. Iteration MAXQUAD with $k=200,L=1000$ in Table 2.
@@
  • rAPEX (Algorithm 6): implemented with Julia, the subproblem is solved by NonNegLeastSquares.jl.
  • restarted APL [44]: implemented with Julia, the lower bound oracle is solved by Gurobi and the upper bound oracle is solved by NonNegLeastSquares.jl.
  • Proximal Bundle Method [37]: implemented with Julia, the subproblem is solved by Ipopt.
  • Subgradient regularized descent (SRD) and SRD_adapt [49]: implemented with Matlab, see website.
@@ 8.2 Two-stage stochastic linear programming
$\min_{l_{x}\leq x\leq u_{x},x\in\mathbb{R}^{n_{1}}}\ c_{1}^{\top}x+\mathbb{E}_{\xi\sim\Xi}\left[Q\left(x;\xi\right)\right]\ \operatornamewithlimits{s.t.}\ A_{1}x?b_{1}\ ,\ A_{1}\in\mathbb{R}^{m_{1}\times n_{1}}$ (62)
@@
$\begin{aligned} Q\left(x;\xi\right)=\min_{l_{y}\leq y\leq u_{y},y\in\mathbb{R}^{n_{2}}}&c\left(\xi\right)^{\top}y\ \\ \operatornamewithlimits{s.t.}\ &A_{1}\left(\xi\right)x+A_{2}\left(\xi\right)y?b_{2}\left(\xi\right),A_{1}(\xi)\in\mathbb{R}^{m_{2}\times n_{1}},A_{2}(\xi)\in\mathbb{R}^{m_{2}\times n_{2}}\end{aligned}.$ (63)
@@
$T_{\varepsilon}\left(\mathcal{A},f_{U}\right)>d\ \text{or\ }T_{\varepsilon}\left(\mathcal{A},f_{U}\right)=T_{\varepsilon}\left(\mathcal{Z}_{\mathcal{A}},f\right)\ .$ (64)
@@
$z^{t}=U^{\top}a^{t}\ \text{and}\ \operatorname{supp}\left\{z^{t}\right\}\subseteq S_{t}\ ,\forall t\leq d\ .$ (65)
In the following, we begin to prove the equivalent claim (65). We explicitly construct algorithm $\mathcal{Z}_{\mathcal{A}}$ and matrix $U$ to meet the requirement in (65). The basic idea is that for any subsequence $\left\{a^{t}\right\}_{t\leq d}$, there exists a matrix set $\mathcal{U}\left(t\right)$ and algorithm class $\mathbb{A}\left(t\right)$ such that every matrix $U\in\mathcal{U}\left(t\right)$ and every algorithm $\mathcal{Z}_{\mathcal{A}}\in\mathbb{A}\left(t\right)$ satisfies our requirement. Furthermore, we can show that the intersection of these sets are nonempty, hence, we can always choose the element in the intersection to meet our requirement in (65). For the base case, we define the matrix set and function class set as
$\begin{split}\mathcal{U}\left(1\right):=&\left\{U\in\mathbb{R}^{d^{\prime}\times d}:U^{\top}U=I_{d},\left\langle u_{i},v_{1}\right\rangle=0,\forall i\in\left[d\right]\right\}\\ \mathbb{A}\left(1\right):=&\left\{\mathcal{Z}:\mathcal{Z}\left(\varnothing\right)=0\right\}\end{split}$ (66)
@@
$\begin{split}&U\in\mathcal{U}\left(t\right)=\left\{U\in\mathbb{R}^{d\times d}:U^{\top}U=I_{d},\left\langle u_{i},v_{j}\right\rangle=0,\forall i\in\left[d\right]\backslash S_{t},\forall j\leq t\right\}\cap\mathcal{U}\left(t-1\right)\\ &\mathcal{Z}_{\mathcal{A}}\in\mathbb{A}\left(t\right)=\left\{\mathcal{Z}:\begin{array}[]{c}\mathcal{Z}\left(\nabla f\left(z^{1}\right),\ldots,\nabla f\left(z^{t-1}\right)\right)=U^{\top}\mathcal{A}\left(U\nabla f\left(z^{1}\right),\ldots,U\nabla f\left(z^{t-1}\right)\right),\\ U\in\mathcal{U}\left(t\right),\mathcal{Z}\in\mathbb{A}\left(t-1\right)\end{array}\right\},\forall t\geq 2\end{split}$ (67)
Before proceeding, we first give more detailed illustration about the construction of the two set $\mathcal{U}\left(t\right)$ and $\mathbb{A}\left(t\right)$. The construction of $\mathbb{A}\left(t\right)$ is to meet the first requirement of (65). To keep the relation between that $z^{t}=U^{\top}a^{t}$ holds for any $t$, the algorithm $\mathcal{Z}_{\mathcal{A}}$ must be use the information of initial algorithm $\mathcal{A}$ and try to emulated the output sequence that generated by $\mathcal{A}$. Hence, we select the direct mapping that $\mathcal{Z}_{\mathcal{A}}\left(\nabla\right)=U^{\top}\mathcal{A}\left(U\nabla\right)$. Here we use notation $\nabla$ to denote the information of gradient. The construction of $\mathcal{U}\left(t\right)$ is aim to meet the second requirement of (65). Once $z^{t}=U^{\top}a^{t}$ holds, then $\operatorname{supp}\left\{z^{t}\right\}\subseteq S_{t}$ holds only depends on $u_{i},\forall i\in\left[d\right]\backslash S_{t}$ is located in the complementary of $\operatorname{Span}\left\{v_{1},\ldots,v_{t}\right\}$.
Next, we show that every matrix $U\in\mathcal{U}\left(t\right)$ and every algorithm $\mathcal{Z}_{\mathcal{A}}\in\mathbb{A}\left(t\right)$ satisfies our requirement (65) with $t$. Let us prove the argument (65) with the inductive. Since $a^{1}$ is generated by algorithm $\mathcal{A}$ from scratch, i.e., $a^{1}=\mathcal{A}\left(\varnothing\right)$. Hence, $a^{1}\in\mathbb{R}^{d}$ is an arbitrary (but deterministic) vector. We thus satisfy (65) by selecting $U\in\mathcal{U}\left(1\right)$ and $z^{1}=0$, i.e., $0=\mathcal{Z}_{\mathcal{A}}\left(\varnothing\right)$. Obviously, $\mathcal{U}\left(1\right)\neq\varnothing$. Hence, at iteration $t=1$, we can find proper matrix $U$ and $\mathcal{Z}_{\mathcal{A}}$ satisfy (65). Suppose construction of $\mathcal{Z}_{\mathcal{A}}\in\mathbb{A}\left(t\right)$ and $U\in\mathcal{U}\left(t\right)$ in (67) such that (65) holds for any $s<t$, i.e., $z^{s}=U^{\top}a^{s},$ $\operatorname{supp}\left\{z^{s}\right\}\subseteq S_{s}$ and $\mathcal{U}\left(s\right)\neq\varnothing$ for all $s<t\leq T_{0}$. Now, we show that $z^{t}=U^{\top}a^{t}$ and $\operatorname{supp}\left\{z^{t}\right\}\subseteq S_{t}$ hold.
@@
$z^{t}=\mathcal{Z}_{\mathcal{A}}\left(\nabla f\left(z^{1}\right),\ldots,\nabla f\left(z^{t-1}\right)\right)=U^{\top}\mathcal{A}\left(U\nabla f\left(z^{1}\right),\ldots,U\nabla f\left(z^{t-1}\right)\right)=U^{\top}a^{t}$ (68)

### v2

@@ 6 Restarted APEX for unknown $\mu^{*}$
By Lemma 6.2, Algorithm 6 will generate a point set $\mathcal{P}_{s}$ as $\left(\sqrt{2\Delta_{s}/\mu_{s}},\sqrt{\mu_{s}\Delta_{s}/2}\right)$-W-certificate for point $\bar{y}^{s-1}$ in function $f$. By Lemma 4.2, the W-certificate and $\mu^{*}$-strong convexity implies that $f\left(\bar{y}^{s-1}\right)-f^{*}\leq\max\left\{1,\mu_{s}/\mu^{*}\right\}\Delta_{s}$. To make sure $f\left(\bar{y}^{s-1}\right)-f^{*}\leq\varepsilon$, we need $\mu_{s}\Delta_{s}\leq\min\left\{\mu_{s},\mu^{*}\right\}\varepsilon$. Notice that Theorem 6.1 is to bound the iteration number for condition $\mu_{s}\Delta_{s}\leq 2\varepsilon^{2}$. Hence, replacing $2\varepsilon^{2}$ in the proof procedure of Theorem 6.1 with $\min\left\{\mu_{s},\mu^{*}\right\}\varepsilon$} and noting that ${\mu}_{s}\geq\frac{1}{4}\mu^{*}$, we have the complexity result. ∎
@@ 7 Lower oracle complexity bound for $\mu$-QG and $\left(k,L\right)$-PWS optimization
The proof follows the two-stage approach pioneered by [11, 12]. First, we show that the lower complexity of any first-order deterministic algorithm (Definition 6) is lower bounded by that of any zero-respecting algorithm (Definition 7) in Proposition 7.1. It thus suffices to prove the lower bound within the zero-respecting algorithm class. Second, inspired by Nesterov’s construction [65], we exhibit a simple $\mu$-strongly convex, $\left(k,L\right)$-PWS function for which every zero-respecting method indeed incurs $\Omega\left(k\sqrt{{L}/{\mu}}\log\left({1}/{\varepsilon}\right)\right)$ queries. Since strongly convex functions are also $\mu$-QG functions, this result also applies to $\mu$-QG functions.
We begin by introducing precise definitions under consideration, including function class (Definition 4), orthogonal connection matrix (Definition 5), first-order deterministic algorithm (Definition 6) and first-order zero-respecting sequences and algorithm (Definition 7).
@@
Definition 5 (Orthogonal connection matrix).
@@
Definition 6 (First-order deterministic algorithm).
@@
$x^{i}=\mathcal{A}\left(\nabla f\left(x^{1}\right),\ldots,\nabla f\left(x^{i-1}\right)\right)\ ,\ \forall i\in\mathbb{N}\ .$ (7.1)
Definition 7 (First-order zero-respecting algorithm).
@@
$\operatorname{supp}\left\{x^{t}\right\}\subseteq\cup_{s<t}\operatorname{supp}\left\{\nabla f\left(x^{s}\right)\right\}\ ,\ \text{for each }t\in\mathbb{N}.$ (7.2)
@@
$T_{\varepsilon}\left(\mathcal{A}\left[f\right],f\right):=\inf\left\{t\in\mathbb{N}:f\left(x^{t}\right)-f^{*}\leq\varepsilon\right\}\ .$ (7.3)
@@
$\mathcal{T}_{\varepsilon}\left(\mathbb{A},\mathcal{F}^{d}\right):=\inf_{\mathcal{A}\in\mathbb{A}}\sup_{f\in\mathcal{F}^{d}}T_{\varepsilon}\left(\mathcal{A}\left[f\right],f\right)\ .$ (7.4)
@@
$\phi_{\mu,L}\left(x\right)=\frac{\mu}{2}\left\|x\right\|^{2}+\frac{L-\mu}{8}x^{\top}Ax-\frac{L-\mu}{4}e_{1}^{\top}x\ ,$ (7.5)
@@
$\phi_{\mu,L}^{k}\left(x\right)=\max_{i\in\left[k\right]}\left\{\phi_{\mu,L}\left(x_{(i)}\right)\right\}\ ,\ x=\left[x_{(1,:)}^{\top},x_{(2,:)}^{\top},\ldots,x_{(k,:)}^{\top}\right]\in\mathbb{R}^{d}\ ,\ x_{(i,:)}\in\mathbb{R}^{n_{i}}\ ,\ \forall i\in\left[k\right]\ .$ (7.6)
@@
$\left\|x^{t}-x^{*}\right\|^{2}$ $\geq\left(\frac{\sqrt{L/\mu}-1}{\sqrt{L/\mu}+1}\right)^{2t/k}\left\|x^{0}-x^{*}\right\|^{2}\ ,$ (7.7)
$f\left(x^{t}\right)-f\left(x^{*}\right)$ $\geq\frac{\mu}{2}\left(\frac{\sqrt{L/\mu}-1}{\sqrt{L/\mu}+1}\right)^{2t/k}\left\|x^{0}-x^{*}\right\|^{2}\ ,$ (7.8)
@@
It suffices to prove that the claims (7.7) and (7.8) hold for any first-order zero-respecting algorithm $\mathcal{A}_{zr}$ by Proposition 7.1. We test $\mathcal{A}_{zr}$ on the PWS instance $f\left(x\right)=\phi_{\mu,L}^{k}\left(x\right)$, where in the definition (7.6) we set $n_{i}=\infty,i=1,\ldots,k$. For notation simplicity, we denote $q=\frac{\sqrt{L/\mu}-1}{\sqrt{L/\mu}+1}$. Hence, by the fourth property (Item 4) of $\phi_{\mu,L}\left(x\right)$ that $x^{*}=\operatornamewithlimits{argmin}_{x\in\mathbb{R}^{\infty}}f\left(x\right)$ has a block-coordinate structure that is $x^{*}=\left[\left(x_{(1,:)}^{*}\right)^{\top},\left(x_{(2,:)}^{*}\right)^{\top},\ldots,\left(x_{(k,:)}^{*}\right)^{\top}\right]$, where $x_{(j,i)}^{*}=q^{i}$. Without loss of generality, we assume algorithm starts from $x^{0}=0$. Then by construction, we have
$\left\|x^{0}-x^{*}\right\|^{2}=\sum_{j=1}^{k}\sum_{i=1}^{n_{j}}x_{(j,i)}^{*}=\frac{kq^{2}}{1-q^{2}}\ .$ (7.9)
@@
$\begin{split}&\left\|x^{t}-x^{*}\right\|^{2}\geq\sum_{i=1}^{k}\left\|x_{(i,:)}^{t}-x_{(i,:)}^{*}\right\|^{2}\overset{(a)}{\geq}\sum_{i=1}^{k}\left\|x_{(i,t_{i}+1:\infty)}^{*}\right\|^{2}\\ &\overset{(b)}{=}\sum_{i=1}^{k}\frac{q^{2(t_{i}+1)}}{1-q^{2}}\overset{(c)}{=}\left\|x^{0}-x^{*}\right\|^{2}\sum_{i=1}^{k}\frac{q^{2t_{i}}}{k}\overset{(d)}{\geq}\left\|x^{0}-x^{*}\right\|^{2}q^{2t/k}\end{split}\ ,$ (7.10)
In the above, $(a)$ holds by discarding the first difference norm between the first $t_{i}$ elements of $x_{(i,:)}^{t}$ and $x_{(i,:)}^{*}$ , $(b)$ follows from $x_{(j,i)}^{*}=q^{i}$, $(c)$ holds by (7.9) and $(d)$ is because $g:t\to q^{2t}$ is convex, then $\frac{1}{k}\sum_{i=1}^{k}g\left(t_{i}\right)\geq g\left(\frac{1}{k}\sum_{i=1}^{k}t_{i}\right)=q^{2\bar{t}}$. Hence, we complete the first result (7.7). Combining it and $f\left(x\right)$ is $\mu$-strongly convex, we complete the proof of the second result (7.8). ∎
@@ 8.1 Finite max of convex quadratic functions (MAXQUAD)
Experiment setting. We first evaluate the performance of rAPEX on small-scale MAXQUAD with varying piece number $k$, smoothness $L$ and initial QG estimates $\hat{\mu}_{1}$, as reported in Table 2. In all experiments, we set the dimension ($d$) to $300$, cut number ($m$) to $50$, the optimal QG modulus ($\mu^{*}$) to $1.0$, and terminate rAPEX once rAPEX improves the upper bound (Algorithm 6 enters Line 16) with $\Delta_{s}\leq 10^{-6}$. Ipopt [71] is used as a baseline to provide optimal objective $(f^{*})$.
@@
Figure 5: Absolute error of upper bound and lower bound of rAPEX v.s. Iteration on MAXQUAD with $k=50,L=10$ in Table 2.
Figure 6: Absolute error of upper bound and lower bound of rAPEX v.s. Iteration on MAXQUAD with $k=200,L=1000$ in Table 2.
@@ 8.2 Two-stage stochastic linear programming
$\min_{l_{x}\leq x\leq u_{x},x\in\mathbb{R}^{n_{1}}}\ c_{1}^{\top}x+\mathbb{E}_{\xi\sim\Xi}\left[Q\left(x;\xi\right)\right]\ \operatornamewithlimits{s.t.}\ A_{1}x?b_{1}\ ,\ A_{1}\in\mathbb{R}^{m_{1}\times n_{1}}$ (8.1)
@@
$\begin{aligned} Q\left(x;\xi\right)=\min_{l_{y}\leq y\leq u_{y},y\in\mathbb{R}^{n_{2}}}&c\left(\xi\right)^{\top}y\ \\ \operatornamewithlimits{s.t.}\ &A_{1}\left(\xi\right)x+A_{2}\left(\xi\right)y?b_{2}\left(\xi\right),A_{1}(\xi)\in\mathbb{R}^{m_{2}\times n_{1}},A_{2}(\xi)\in\mathbb{R}^{m_{2}\times n_{2}}\end{aligned}.$ (8.2)
@@
$T_{\varepsilon}\left(\mathcal{A},f_{U}\right)>d\ \text{or\ }T_{\varepsilon}\left(\mathcal{A},f_{U}\right)=T_{\varepsilon}\left(\mathcal{Z}_{\mathcal{A}},f\right)\ .$ (A.1)
@@
$z^{t}=U^{\top}a^{t}\ \text{and}\ \operatorname{supp}\left\{z^{t}\right\}\subseteq S_{t}\ ,\forall t\leq d\ .$ (A.2)
In the following, we begin to prove the equivalent claim (A.2). We explicitly construct algorithm $\mathcal{Z}_{\mathcal{A}}$ and matrix $U$ to meet the requirement in (A.2). The basic idea is that for any subsequence $\left\{a^{t}\right\}_{t\leq d}$, there exists a matrix set $\mathcal{U}\left(t\right)$ and algorithm class $\mathbb{A}\left(t\right)$ such that every matrix $U\in\mathcal{U}\left(t\right)$ and every algorithm $\mathcal{Z}_{\mathcal{A}}\in\mathbb{A}\left(t\right)$ satisfies our requirement. Furthermore, we can show that the intersection of these sets are nonempty, hence, we can always choose the element in the intersection to meet our requirement in (A.2). For the base case, we define the matrix set and function class set as
$\begin{split}\mathcal{U}\left(1\right):=&\left\{U\in\mathbb{R}^{d^{\prime}\times d}:U^{\top}U=I_{d},\left\langle u_{i},v_{1}\right\rangle=0,\forall i\in\left[d\right]\right\}\\ \mathbb{A}\left(1\right):=&\left\{\mathcal{Z}:\mathcal{Z}\left(\varnothing\right)=0\right\}\end{split}$ (A.3)
@@
$\begin{split}&U\in\mathcal{U}\left(t\right)=\left\{U\in\mathbb{R}^{d\times d}:U^{\top}U=I_{d},\left\langle u_{i},v_{j}\right\rangle=0,\forall i\in\left[d\right]\backslash S_{t},\forall j\leq t\right\}\cap\mathcal{U}\left(t-1\right)\\ &\mathcal{Z}_{\mathcal{A}}\in\mathbb{A}\left(t\right)=\left\{\mathcal{Z}:\begin{array}[]{c}\mathcal{Z}\left(\nabla f\left(z^{1}\right),\ldots,\nabla f\left(z^{t-1}\right)\right)=U^{\top}\mathcal{A}\left(U\nabla f\left(z^{1}\right),\ldots,U\nabla f\left(z^{t-1}\right)\right),\\ U\in\mathcal{U}\left(t\right),\mathcal{Z}\in\mathbb{A}\left(t-1\right)\end{array}\right\},\forall t\geq 2\end{split}$ (A.4)
Before proceeding, we first give more detailed illustration about the construction of the two set $\mathcal{U}\left(t\right)$ and $\mathbb{A}\left(t\right)$. The construction of $\mathbb{A}\left(t\right)$ is to meet the first requirement of (A.2). To keep the relation between that $z^{t}=U^{\top}a^{t}$ holds for any $t$, the algorithm $\mathcal{Z}_{\mathcal{A}}$ must be use the information of initial algorithm $\mathcal{A}$ and try to emulated the output sequence that generated by $\mathcal{A}$. Hence, we select the direct mapping that $\mathcal{Z}_{\mathcal{A}}\left(\nabla\right)=U^{\top}\mathcal{A}\left(U\nabla\right)$. Here we use notation $\nabla$ to denote the information of gradient. The construction of $\mathcal{U}\left(t\right)$ is aim to meet the second requirement of (A.2). Once $z^{t}=U^{\top}a^{t}$ holds, then $\operatorname{supp}\left\{z^{t}\right\}\subseteq S_{t}$ holds only depends on $u_{i},\forall i\in\left[d\right]\backslash S_{t}$ is located in the complementary of $\operatorname{Span}\left\{v_{1},\ldots,v_{t}\right\}$.
Next, we show that every matrix $U\in\mathcal{U}\left(t\right)$ and every algorithm $\mathcal{Z}_{\mathcal{A}}\in\mathbb{A}\left(t\right)$ satisfies our requirement (A.2) with $t$. Let us prove the argument (A.2) with the inductive. Since $a^{1}$ is generated by algorithm $\mathcal{A}$ from scratch, i.e., $a^{1}=\mathcal{A}\left(\varnothing\right)$. Hence, $a^{1}\in\mathbb{R}^{d}$ is an arbitrary (but deterministic) vector. We thus satisfy (A.2) by selecting $U\in\mathcal{U}\left(1\right)$ and $z^{1}=0$, i.e., $0=\mathcal{Z}_{\mathcal{A}}\left(\varnothing\right)$. Obviously, $\mathcal{U}\left(1\right)\neq\varnothing$. Hence, at iteration $t=1$, we can find proper matrix $U$ and $\mathcal{Z}_{\mathcal{A}}$ satisfy (A.2). Suppose construction of $\mathcal{Z}_{\mathcal{A}}\in\mathbb{A}\left(t\right)$ and $U\in\mathcal{U}\left(t\right)$ in (A.4) such that (A.2) holds for any $s<t$, i.e., $z^{s}=U^{\top}a^{s},$ $\operatorname{supp}\left\{z^{s}\right\}\subseteq S_{s}$ and $\mathcal{U}\left(s\right)\neq\varnothing$ for all $s<t\leq T_{0}$. Now, we show that $z^{t}=U^{\top}a^{t}$ and $\operatorname{supp}\left\{z^{t}\right\}\subseteq S_{t}$ hold.
@@
$z^{t}=\mathcal{Z}_{\mathcal{A}}\left(\nabla f\left(z^{1}\right),\ldots,\nabla f\left(z^{t-1}\right)\right)=U^{\top}\mathcal{A}\left(U\nabla f\left(z^{1}\right),\ldots,U\nabla f\left(z^{t-1}\right)\right)=U^{\top}a^{t}$ (A.5)